### v2

# Two-phase validation sampling via principal components to improve efficiency in multi-model estimation from error-prone biomedical databases

Abstract
Two-phase sampling offers a cost-effective way to validate error-prone covariate measurements in biomedical databases. Inexpensive or easy-to-obtain information is collected for the entire study in Phase I. Then, a subset of patients undergoes cost-intensive validation (e.g., expert chart review) to collect more accurate data in Phase II. Critically, any Phase I variables can be used to strategically select Phase II, often enriched for a particular model of interest, and then missing data methods like multiple imputation can be used for estimation. However, when balancing primary and secondary analyses, competing models and priorities can result in poorly defined objectives for the most informative Phase II sampling criterion. Extreme tail sampling (ETS), wherein patients with the smallest and largest values of a particular quantity (like a covariate or residual) are selected, can offer great statistical efficiency in two-phase studies when focusing on a single analytic objective by targeting observations with the biggest contributions to the Fisher information. We propose an intuitive, easy-to-use approach that extends ETS to balance and prioritize explaining the largest amount of variability across multiple models of interest. Using principal components, we succinctly summarize the inherent variability of all models’ error-prone exposures. Then, we sample patients with the most extreme principal components for validation. Through simulations and an application to the National Health and Nutrition Examination Survey (NHANES), the proposed sampling strategy offered simultaneous efficiency gains across multiple models of interest. Its advantages persisted across various real-world scenarios: different covariance structures, measurement error severities, validation proportions, and shared versus distinct model outcomes. When designing a validation study, concentrating on a single model may be short-sighted. Allocating resources more broadly, while still strategically, could balance multiple analytical goals simultaneously. Employing dimension reduction before sampling will allow this strategy to scale up well to big-data applications with many error-prone covariates.
Keywords data audit $\cdot$ extreme tail sampling $\cdot$ measurement error $\cdot$ NHANES $\cdot$ principal component analysis
1 Introduction
Data are rarely collected for a single purpose in research. Observational databases, like electronic health records (EHR), compile troves of information on extensive samples, providing new opportunities to explore various novel scientific interests with a single data source. Randomized experiments, such as clinical trials, attempt to simultaneously address the primary endpoint alongside multiple secondary objectives. Moreover, data from both observational and randomized studies require validation in their own ways.
1.1 Necessity of Validation in Observational Studies and Clinical Trials
Observational data are generally routinely collected with other purposes (e.g., providing good clinical care) at the forefront. They may have errors resulting from various sources, like improper collection, transcription, or data cleaning. Thus, when repurposed for research [28], these data may exhibit measurement error [14] and bias downstream analyses [13]. Manual chart reviews, wherein clinical experts compare research data to clinical source documents (like patients’ electronic charts), can validate observational datasets (e.g., [9, 19, 21, 34]), but they are usually implausible for entire studies. Partial validation (i.e., of a subset of patients) is more feasible.
Source document verification (data validation) has long been standard in clinical trials [42], making data collected therein of generally high quality in terms of typos and other human errors. Further, under random assignment, the exposure of interest (generally a treatment) should be perfectly collected. Still, other variables, like biomarkers, may be imperfectly measured. Validation in the form of re-measurement using more accurate instruments can account for errors [15, 22, 35], but the more accurate measures are usually invasive or expensive, making them unreasonable to obtain for all trial participants.
1.2 Informative Validation Sampling Strategies
Partial validation studies are a special type of two-phase design [43]. In Phase I, inexpensive or readily available information is collected for all patients in the study. In Phase II, a subset of patients is chosen for validation, such that more accurate measurements are collected. Sampling for Phase II can leverage any information available for all patients in Phase I, opening the door to many targeted strategies.
Often, the Phase II subset is enriched for a particular model (e.g., to minimize the variance of a model coefficient of interest). This enrichment depends on the model type and estimation procedure. Essentially, identifing which patients would be most informative to validate depends on what we plan to analyze about them and how.
For design-based estimators, like inverse probability weighting (IPW) and mean score, multiple informative sampling designs exist to minimize the variances of a particular model coefficient. Model-based estimators, like multiple imputation and maximum likelihood estimation (MLE), allow for more design flexibility than design-based estimators, since some patients can be excluded from consideration (i.e., given zero validation sampling probability). While the variances of model-based estimators do not lend themselves to intuitive, analytical solutions for optimal designs (as with design-based ones), many have been proposed (e.g., [3, 20, 24, 37]).
When no covariate information is available pre-validation, sampling patients with extreme outcome values (called outcome dependent or extreme tail sampling [ETS]) is typically more efficient than simple random sampling (SRS) [1, 2, 11, 17, 29, 31, 45]. When covariate information is available, other ETS strategies, including sampling based on (weighted) residuals [8, 21, 23, 37] and influence functions [3], can offer sizable efficiency gains, as well.
1.3 Balancing Multiple Models and Objectives
Most informative two-phase sampling designs prioritize estimating a single coefficient well (e.g., with great precision). Yet, post-hoc and exploratory analyses can result in competing objectives when designing a validation study. Balancing multiple analytical goals across various models leads to an ill-defined optimization problem, which complicates the optimal approach for validation sampling. Of course, a “silver bullet” would maximize all objectives (unrealistic in practice), whereas maximizing just one leads to a sub-optimal system solution. For instance, focusing on a single objective may give us good performance for that model at the detriment of all others. Various approaches exist to the multi-objective problem [7], but, to our knowledge, none have been applied for two-phase designs.
Optimally designing a validation study to efficiently estimate multiple coefficients in the same model poses a related, yet unique challenge. This setting, wherein a single model is of interest and high statistical precision for many or all coefficients is desired, has received more attention in the two-phase literature than ours with multiple models. For example, Sauer et al. (2021) [31] demonstrate the trade-off between focusing a two-phase design strategy on one versus many model coefficients. Considering optimal allocation strategies that minimize (i) the variance of a single coefficient and (ii) the trace of the coefficients’ variance-covariance matrix (i.e., the sum of their variances), they show that while the latter offers smaller efficiency gains for any particular coefficient it offers better precision for the model as a whole.
Intuitively, we expect the same trade-off when estimating coefficients across multiple models: We can be great for one coefficient and poor for the rest, or we can be good for all of them. While extending existing optimal designs from one model to many is possible, many depend upon functions of the moments of the coefficients of interest (often unknowable in practice). Herein, we propose a simpler approach that scales up well with the number of models and variables, making it particularly appealing for big-data applications.
Principal component analysis (PCA), a popular nonparametric dimension reduction technique, offers a promising way to adapt existing informative sampling designs to accommodate multiple modeling objectives. The principal components succinctly summarize the variability within the data, and, conveniently, applying PCA on centered and scaled data or on the data’s covariance matrix inherently resolves the issue of differences in the variables’ scales. We explore how PCA can capture information across multiple error-prone variables, each of interest in its own statistical model, and guide targeted sampling strategies for partial validation.
1.4 National Health and Nutrition Examination Survey
We consider the accuracy of food recall logs and dietary measurements in the National Health and Nutrition Examination Survey (NHANES) [12]. For the 2021–2023 cohort, two dietary recalls were conducted via phone interviews. Individuals reported the foods and drinks they had consumed within the past $24$ hours.
While convenient, even the gold standard method of food recall logs is subject to recall bias, under- or over-reporting, and fallibility of portion size estimation (see [38] and references therein). Converting food intake into nutrients can be further complicated by product differences in nutrient density when individuals oversimplify food items (e.g., the amount of protein in “yogurt”). We estimate the relationship between error-prone nutrient intake and relevant health outcomes, comparing statistical precision under various sampling strategies.
1.5 Overview
We propose a new informative sampling strategy that balances multiple modeling objectives simultaneously. Validation sampling depends on the first principal component of all observed error-prone exposures, and the outcome variables for the multiple models may be shared or unique. Our design is simple to implement and balances competing analytical objectives in an intuitive and easy-to-understand approach. The resulting partially validated data are analyzed using multiple imputation, but other model-based methods could be used. In simulations and an application to error-prone, self-reported food intake data from the NHANES study, we show that the proposed design outperforms both SRS and ETS focused on a single model in terms of statistical efficiency (i.e., the inverse of the estimated coefficients’ variances).
Statement of Significance
  • Problem Two-phase sampling offers a cost-effective way to partially validate error-prone measurements in biomedical databases (e.g., through expert chart review or evaluation with more precise instruments).
  • What Is Already Known Any fully-observed variables can be used to strategically select the validation subset, which is often enriched for a particular model of interest, and then missing data methods like multiple imputation can be used for estimation from partially validated data. However, nearly all such targeted validation sampling strategies promote efficiency in one or more coefficients in a single model.
  • What This Paper Adds We propose an intuitive, easy-to-use approach that balances and prioritizes explaining the largest amount of variability across multiple models of interest. First, principal components analysis is used to summarize variability in the error-prone data, and then observations with the most extreme principal components (i.e., the largest or smallest) are chosen for validation.
2 Methods
2.1 Model and Data
Suppose we are conducting a study of $N$ patients on whom $J$ separate statistical models are of interest. For each model $j$, we have a continuous outcome $Y_{ji}$, an exposure of interest $X_{ji}$, and a vector of other confounders ${\mathbf{Z}}_{ji}$ ($j\in\{1,\dots,J\}$, $i\in\{1,\ldots,N\}$). These variables are assumed to be related through the usual normal linear regression model:
$Y_{ji}=\beta_{0j}+\beta_{1j}X_{ji}+{\boldsymbol{\beta}}_{2j}{\mathbf{Z}}_{ji}+\epsilon_{ji},$ (1)
where the random errors $\epsilon_{ji}$ are independent and identically distributed (i.i.d.) following a normal distribution with mean =0=0 and variance $\sigma_{j}^{2}$. We are most interested in estimating $\beta_{1j}$ from each model, but all the exposures $\boldsymbol{X}_{i}=(X_{1i},\dots,X_{Ji})^{\top}$ are difficult to measure accurately.
Instead, we observe error-prone versions $\boldsymbol{X}^{*}_{i}=\boldsymbol{X}_{i}+\boldsymbol{U}_{i}$ for all $N$ patients, where $\boldsymbol{U}_{i}=(U_{1i},\dots,U_{Ji})^{\top}$ is the vector of additive measurement errors for patient $i$’s exposures from all $J$ models. We are planning a partial validation study, such that a subset of $n$ patients ($n<N$) will have $\boldsymbol{X}_{i}$ measured. To satisfy the missing at random (MAR) assumption [18] for our ultimate analyses, validation sampling can depend only on information that is fully observed (i.e., the outcomes, error-prone exposures, and confounders). Without loss of generality, we will assume that if patient $i$ is sampled for validation, then $\boldsymbol{X}_{i}$ is fully observed. We want to use these partially validated data to estimate the $J$ separate models, all of which have equal priority in this study.
Throughout, our priority is to strategically design the validation study to promote the efficiency of the coefficient estimates $\hat{{\boldsymbol{\beta}}}_{1}=(\hat{\beta}_{11},\dots,\hat{\beta}_{1J})^{\top}$, accompanying the exposures $\boldsymbol{X}_{i}$, obtained via multiple imputation. Other model-based estimation approaches, like MLE, should benefit from similar efficiency gains under the proposed design. We focused on multiple imputation here for its simpler implementation.
2.2 Validation Sampling Designs
Traditionally, targeted sampling focuses on one model $j$, and the goal is to identify which $n$ patients, if validated, would minimize the variance of $\hat{\beta}_{1j}$ corresponding to the exposure of interest $X_{ji}$ [3, 15, 20, 21, 24, 34, 37, 46]. Designs have also been proposed to improve estimation across multiple coefficients ($\hat{\beta}_{1j},\hat{\boldsymbol{\beta}}_{2j}$) in a given model by minimizing the trace (A-optimality) or determinant (D-optimality) of their variance-covariance matrix [10, 31]. However, these strategies do not consider the estimation of the coefficients from the other J1J-1 models.
Most of these designs are already complex for a single coefficient from one chosen model; extending them to consider multiple coefficients simultaneously, potentially each with its own outcome, would be challenging. Further, theoretically optimal designs to estimate $\hat{\beta}_{1j}$ are functions of true $\beta_{1j}$. Multi-wave strategies to approximate theoretically optimal designs are becoming commonplace [20, 21, 25, 34], but they can be more logistically challenging. Therefore, we instead propose a novel strategy for partial validation in a multi-objective study based on the ETS design (already a common strategy for single-objective studies) and using PCA to create a single measure summarizing variability across all models’ error-prone exposures.
In exploratory analyses or feature creation, PCA is a common approach to dimension reduction. Through a smaller subset of newly created variables (called principal components or loadings), PCA seeks to explain variability in the original data. This property is central to our use of PCA in validation study design. We apply PCA to all the models’ error-prone exposures $\boldsymbol{X}^{*}$ (and, if desired, additional confounders) to create a single variable capturing the variability between them. In particular, we take the first principal component $PC_{1}^{*}$, which will always explain the most variability out of all the estimated principal components. Then, we prioritize patients for validation based on $PC_{1}^{*}$, rather than focusing only on a single $X_{j}^{*}$ or trying to balance all of $\boldsymbol{X}^{*}$ simultaneously. This strategy balances the reduction in variance across all $J$ models without adding computational burden, and it seamlessly integrates with existing single-objective validation sampling methods, like ETS.
Throughout this paper, we consider three sampling strategies for the validation of $n$ patients out of a study of size $N$. As the baseline design, we include simple random sampling (SRS). The $n$ patients are chosen with equal probability ($n/N$), which ignores all available information and makes $\boldsymbol{X}$ missing completely at random (MCAR) [18]. To represent a typical existing strategy, wherein the validation study is enriched for the model of primary interest, we include extreme tail sampling on the primary model’s error-prone exposure (ETS-$X_{p}^{*}$). The original $N$ patients are ordered ascendingly by $X_{p}^{*}$ ($p\in\{1,\dots,J\}$), and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation. In practice, the primary model would be chosen a priori, for example, based on clinical interest. Finally, we implement our proposed strategy, extreme tail sampling on the first principal component of all error-prone exposures (ETS-$PC_{1}^{*}$). The original $N$ patients are ordered ascendingly by $PC_{1}^{*}$, the first principal component derived from $\boldsymbol{X}^{*}$ for all $J$ models, and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation.
2.3 Intuition Behind Extreme Tail Sampling
The intuition behind ETS designs comes from the belief that extreme observations (i.e., those in the tails of the distribution) are most informative to conducting statistically efficient analyses. The statistical underpinnings behind this intuition stem from the large-sample theory behind MLE (the method by which common statistical models are fit).
2.3.1 For the Primary Model
Recall that the primary model of interest captures the relationship between $Y_{p}$ and $X_{p}$ given ${\mathbf{Z}}_{p}$. In linear regression, observations with covariate values $X_{p}$ that are farthest from the mean $\overline{X}_{p}$ make the largest contributions to the Fisher information for the corresponding model coefficient $\beta_{1p}$ and help to reduce variability in its estimate $\hat{\beta}_{1p}$. Thus, these extreme values of $X_{p}$ are deemed “most informative” when estimating our model, and, if can only measure the exposure for some subset of $n$ patients, focusing on those the smallest and largest values should offer the best statistical precision. (This fourth implausible strategy could be denoted by ETS-$X_{p}$.)
However, we cannot prioritize patients for validation based on their values of $X_{p}$ (which is attainable only through validation), and the ETS-$X_{p}^{*}$ design is perhaps the most natural alternative. The hope is that by sampling on the extremes of the error-prone version $X_{p}^{*}$, we can capture many of the same patients as we would have under the ETS-$X_{p}$ design. The efficiency advantages of the ETS-$X_{p}^{*}$ relative to the ETS-$X_{p}$ will depend on how informative $X_{p}^{*}$ is about $X_{p}$. See Figure 1 for an illustration, wherein we compare which patients would be selected for validation based on sampling from the extremes of the true versus error-prone exposure under increasing error severities.
Figure 1: Plot of the primary model’s true exposure $X_{p}$ against the error-prone version $X_{p}^{*}=X_{p}+U$ under varied error severities, driven by the variance of the errors $U$. Each plot contains $N=1000$ simulated observations, and the points are colored by their validation status based on extreme tail sampling on $X_{p}^{*}$ (ETS-$X_{p}^{*}$) versus extreme tail sampling on $X_{p}$ (ETS-$X_{p}$). “Counts” denotes the number of observations in each setting with the corresponding validation status. As the error variance increases, and $X_{p}^{*}$ becomes less informative about $X_{p}$, the number of observations that would be sampled by both ETS designs decreases, and the efficiency of ETS-$X_{p}^{*}$ is expected to drop below that of ETS-$X_{p}$. Still, even under the worst error setting, the ETS-$X_{p}^{*}$ design should capture more information than simple random sampling (SRS).
2.3.2 For Multiple Models
With competing analytic objectives, we shift our design focus to reducing the sum of the variances for coefficients $\hat{\beta}_{1j}$ $(j\in\{1,\ldots,J\})$ across all models. Hereafter, this sum is referred to as the total coefficient variability across all models. Improving the estimation of $\beta_{1p}$ in the primary model, through the ETS-$X_{p}^{*}$ design or otherwise, does not guarantee any improvement in the estimation of coefficients for any other models, and thus it will not help to minimize this total. As before, the most informative observations to each individual model would be those with the most extreme exposures $\boldsymbol{X}$. Still, these values are not available when designing the validation stage, so the next-best thing is the error-prone versions $\boldsymbol{X}^{*}$. However, there is now an added challenge of tailoring the design strategy to prioritize the most extreme values across multiple variables at once.
To this end, we consider how to describe the variability across across all $\boldsymbol{X}^{*}$ in a single dimension and then incorporate that information into validation sampling. Conducting principal component analysis and taking the first principal component, $PC_{1}^{*}$, we obtain a single transformed variable that describes the greatest variability in the set of exposures $\boldsymbol{X}^{*}$, providing a compact criterion on which to sample while balancing all modeling objectives. Our ETS-$PC_{1}^{*}$ design is similarly motivated as the ETS-$X_{p}^{*}$ design. By expressing the entirety of $\boldsymbol{X}^{*}$ in a lower dimension as $PC_{1}^{*}$, we seek to validate patients with the most extreme values of $\boldsymbol{X}^{*}$ (and, ideally, $\boldsymbol{X}$), thereby improving estimation efficiency across all $J$ linear regression models. In Figure 2, we broaden the simulated example from above to evaluate five models of interest.
Figure 2: Plot of all models’ true exposure $X_{j}$ against the first principal component $PC_{1}^{*}$ summarizing all the error-prone exposures $\boldsymbol{X}^{*}$, focusing on the moderate error setting (error variance =0.5=0.5). Each plot contains the same $N=1000$ simulated observations, and the points are colored by their validation status based on extreme tail sampling on the first principal component $PC_{1}^{*}$ (ETS-$PC_{1}^{*}$) versus extreme tail sampling on the specific model’s exposure $X_{j}$ (ETS-$X_{j}$). “Counts” denotes the number of observations in each setting with the corresponding validation status. Across all five models, the number of overlapping observations that would be sampled by both ETS designs is relatively stable, and the efficiency of ETS-$PC_{1}^{*}$ is expected to offer similarly good approximations to the ETS-$X_{j}$ for each model separately.
The ETS-$PC_{1}^{*}$ design’s ability to reduce the total coefficient variability across all models depends on two characteristics of the error-prone exposure data. First, as in the primary model case, the efficiency gains are driven by how much information we can glean from the validated exposures $\boldsymbol{X}$. When the errors in $\boldsymbol{X}^{*}$ are less noisy, sampling on its extremes is expected to help us capture more informative observations in the extremes of $\boldsymbol{X}$. Second, our ability to sample the most extreme observations with respect to all of $\boldsymbol{X}^{*}$ (and subsequently $\boldsymbol{X}$) depends on how much variability is explained by the first principal component $PC_{1}^{*}$. When the error-prone exposures $\boldsymbol{X}^{*}$ are highly correlated, $PC_{1}^{*}$ will be more informative and sampling on its extremes will lead to higher efficiency gains. In the conservative case where the $\boldsymbol{X}^{*}$ are completed uncorrelated, the PCA will yield a separate factor for each exposure, and the ETS-$PC_{1}^{*}$ design will be equivalent to an ETS-$X_{j}^{*}$. In practice, when the exposures are completely independent, nominal correlation between them due to random chance can result in a slight inefficiency when using the proposed design compared to sampling on a single variable.
2.4 Imputation Procedure
In our validated subset, we can estimate the $j$th linear regression model in (1) using the observed exposure $X_{ji}$ directly. For the rest of the patients in the study, we will instead replace their missing value with an informative placeholder $\widehat{X}_{ji}$. Specifically, we adopt a single deterministic imputation framework in the Simulations (Section 3). (Single imputation is computationally quicker than multiple, and we focused on the estimates’ empirical variability to compare the designs.) However, multiple imputation is preferred if estimated variances are desired (e.g., for inference), in which case the imputation models would need to be modified. R code for imputation can be found online at https://github.com/sarahlotspeich/ETS_PCA.
2.4.1 Different Imputation Models Under Different Validation Study Designs
For an unvalidated patient, $\widehat{X}_{ji}$ comes from an imputation model fit to the validation study. Any fully available information that helps estimate $X_{ji}$ can be incorporated, and any variables used in the validation study design must be included [46].
Under an SRS design, the imputed values are defined simply from the error-prone exposure $X_{ji}^{*}$ and the additional confounders ${\mathbf{Z}}_{i}$ as
$\widehat{X}_{ji}=\widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\boldsymbol{\alpha}}_{2j}{\mathbf{Z}}_{i}.$ (2)
When $X_{ji}$ is numeric, the estimated coefficients in (2) can be obtained via normal linear regression using only data from the validation subset. Of course, different regression models could instead be used to impute other types of exposures.
If the validation study was selected under an ETS-$X_{j}^{*}$ design, the error-prone exposure $X_{j}^{*}$ used in the design must be included in all exposures’ imputation models. Let $j^{\prime}$ index which model was used in the design ($j^{\prime}\in\{1,\dots,J\}$). Then, the imputed values $\widehat{X}_{ji}$ are:
$\widehat{X}_{ji}$ $=\begin{cases}\widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\boldsymbol{\alpha}}_{2j}Z_{i}\text{ if }j=j^{\prime}\text{, and }\\ \widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\boldsymbol{\alpha}}_{2j}Z_{i}+\widehat{\alpha}_{3j}X_{ji}^{*}\text{ otherwise.}\end{cases}$ (3)
Note that the coefficients will take on different values in the first versus second case.
Under our proposed ETS-$PC_{1}^{*}$ design, the first principal component $PC_{1}^{*}$ must be included in all imputation models. The imputed values $\widehat{X}_{ji}$ are then defined as
$\widehat{X}_{ji}$ $=\widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\alpha}_{2j}Z_{i}+\widehat{\alpha}_{3j}PC^{*}_{1i}.$ (4)
As with the SRS design, estimating the imputation models under either ETS design depends on the type of covariates, but can be done using standard regression techniques. For ease of notation, $\widehat{\boldsymbol{\alpha}}$ are the estimated coefficients for any of the imputation models.
Single imputation following (2)–(4) has been called deterministic imputation, since missing values are replaced with a single model prediction $\widehat{X}_{ji}$ and treated as fixed in the subsequent analyses [6]. This approach can be implemented using existing software to (i) fit the imputation model and (ii) calculate the predictions. Alternatively, the mice function in the mice R package can be used with argument method = "norm.predict" [41].
2.4.2 Modifications for Multiple Imputation
When using multiple imputation, the imputation models need to be modified one step further. Due to the stochastic nature of multiple imputation, which incorporates variability about the imputation procedure itself, the analysis model outcomes $Y_{j}$ must be included in the imputation models. Otherwise, the imputation models will not be congenial with the analysis model, and the resulting multiple imputation estimates will be biased [6, 26].
There are various ways one might incorporate randomness into the multiple rounds of imputation. One of the earliest and most common methods is to (i) resample the coefficients $\widetilde{\boldsymbol{\alpha}}$ from a distribution based on the fitted ones $\widehat{\boldsymbol{\alpha}}$ in the validation study and (ii) generate the imputed values $\widehat{X}_{ji}$ from the distribution dictated by $\widetilde{\boldsymbol{\alpha}}$ [30]. For example, if $X_{ji}$ is continuous, then in each imputation the parameters $\widetilde{\boldsymbol{\alpha}}$ and $\tilde{\Sigma}$ will first be randomly drawn from a multivariate Normal distribution with mean vector $\widehat{\boldsymbol{\alpha}}$ and variance-covariance matrix $\widehat{\Sigma}$. This process is repeated many times, the analysis model is fit to each imputed dataset, and the estimates from each fitted model are ultimately pooled using Rubin’s rules [30]. We can again use the mice R package. Specifically, the mice function can be used with argument method = "norm", and it will include the outcome $Y_{j}$ in the imputation model by default.
3 Simulations
3.1 Data Generating Mechanism
We generated data for samples of $N=1000$ patients, $n$ of whom had their exposures validated ($n\in\{100,150,250\}$). Each patient had a binary confounder $Z$ generated from a Bernoulli distribution with $\Pr(Z=1)=0.3$. Then, true exposures $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and one of three covariance matrices: independence: $\boldsymbol{\Sigma}_{X}=\boldsymbol{I}_{5}$, equal dependence: $\boldsymbol{\Sigma}_{X}=0.5\boldsymbol{I}_{5}+0.5\boldsymbol{J}_{5}$, or unequal dependence:
$\boldsymbol{\Sigma}_{X}=\begin{bmatrix}1.00&0.05&0.10&0.20&0.35\\ 0.05&1.00&0.15&0.25&0.40\\ 0.10&0.15&1.00&0.30&0.45\\ 0.20&0.25&0.30&1.00&0.50\\ 0.35&0.40&0.45&0.50&1.00\end{bmatrix}$,
where $\mathbf{0}_{5}$ is a $5\times 1$ vector of zeros, $\boldsymbol{I}_{5}$ the $5\times 5$ identity matrix and $\boldsymbol{J}_{5}$ is a $5\times 5$ matrix with ones on the off-diagonal and zeros on the diagonal. Each exposure in $\boldsymbol{X}$ was related to its own numeric response:
$\boldsymbol{Y}$ $=\begin{bmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\\ Y_{4}\\ Y_{5}\end{bmatrix}=\begin{bmatrix}0.5X_{1}+0.1Z\\ 1+X_{2}+0.2Z\\ 2+1.5X_{3}+0.3Z\\ 3+2X_{4}+0.4Z\\ 4+2.5X_{5}+0.5Z\end{bmatrix}+\boldsymbol{\epsilon}$ (5)
for a vector of independent random errors $\boldsymbol{\epsilon}$ generated from a standard multivariate normal distribution. The error-prone versions of these exposures were constructed as $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$, where the measurement errors $\boldsymbol{U}$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\sigma_{U}^{2}\boldsymbol{I}_{5}$ ($\sigma_{U}^{2}\in\{0.06,0.25,1\}$).
3.2 Metrics for Comparison
The SRS, ETS-$X_{1}^{*}$, and ETS-$PC_{1}^{*}$ validation study designs were considered. Without loss of generality, the model of $Y_{1}$ given $(X_{1},Z)$ was treated as primary. When discussing individual models, we focus on comparing the empirical efficiency for the coefficient estimates $\widehat{{\boldsymbol{\beta}}}_{1}$ corresponding to the exposures $\boldsymbol{X}$. When evaluating design performance across all modeling objectives simultaneously, we examine the sum of these coefficients’ empirical variances.
For each of $1000$ simulated datasets per setting, separate linear regression models were estimated following the single imputation approach from Section 2.4. These replicates are summarized by the empirical efficiencies for $\widehat{\beta}_{1j}$ ($j\in\{1,2,3,4,5\}$) in each of the five models (i.e., the inverse of the empirical variances) and the empirical total coefficient variability across all models, $\sum_{j=1}^{5}\widehat{\textrm{V}}(\hat{\beta}_{1j})$.
3.3 Varied Exposures’ Covariance Structure
We explored three different covariance structures for the exposures $\boldsymbol{X}$ described in Section 3.1 and fixed $\sigma^{2}_{U}=1$ and $n=100$. These settings compared the validation study designs under independence versus dependence between the error-prone exposures, since the principal components are expected to be informative only in the latter scenario. Thus, the ETS-$PC_{1}^{*}$ design is expected to perform best when the covariates $\boldsymbol{X}^{*}$ are correlated.
The imputation estimator was empirically unbiased for all models, designs, and covariance structures (Supplemental Figure S1). As expected, the total empirical variability across all models was smallest for the ETS-$PC_{1}^{*}$ design when the exposures $\boldsymbol{X}$ were dependent (either equally or unequally), such that $PC_{1}^{*}$ was informative about $\boldsymbol{X}^{*}$ and, to a lesser extent, $\boldsymbol{X}$ (Figure 3A). Only when the exposures were independent did the ETS-$X_{1}^{*}$ design lead to a slightly smaller total, and ETS-$PC_{1}^{*}$ still beat SRS in that case.
Figure 3: Simulation results comparing the empirical total coefficient variability across all models $\sum_{j=1}^{5}\widehat{\textrm{V}}(\hat{\beta}_{1j})$ under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. In A), three different covariance structures for the five exposures $X_{1},\dots,X_{5}$ were considered. In B), three different variances $\sigma_{U}^{2}$ for the additive measurement errors $U_{1},\dots,U_{5}$ in exposures $X_{1},\dots,X_{5}$ were considered. In C), three different proportions of validated patients out of $N=1000$ were considered.
When broken down by model, there were clear efficiency gains for all models (of $Y_{2},\dots,Y_{5}$) with the ETS-$PC_{1}^{*}$ design (Figure 4A). These gains were comparable in size across all secondary models under equal dependence, and there was hardly any efficiency loss to the ETS-$X_{1}^{*}$ for the primary model. Under unequal dependence, the largest efficiency gain was seen for the model of $Y_{5}$, which had the exposure $X_{5}$ with the highest correlations, on average, to the others. Under unequal covariance, the efficiency for the primary analysis when validating with ETS-$PC_{1}^{*}$ fell farther behind the ETS-$X_{1}^{*}$ design. However, this model-specific loss was offset by lower total empirical variability across all models, as desired. Only when the exposures were independent did the ETS-$PC_{1}^{*}$ design lose to ETS-$X_{1}^{*}$ at both the model-specific and all-model levels, and it still beat SRS.
Figure 4: Simulation results comparing the empirical efficiency under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. In A), three different covariance structures for the five covariates $X_{1},\dots,X_{5}$ were considered. In B), three different variances $\sigma_{U}^{2}$ for the additive measurement errors $U_{1},\dots,U_{5}$ in exposures $X_{1},\dots,X_{5}$ were considered. In C), three different proportions of validated patients out of $N=1000$ were considered.
3.4 Varied Exposures’ Error Severity
To isolate the contribution of the exposure error severity, we changed the variance $\sigma_{U}^{2}\in\{0.25,0.5,1\}$ used to generate the errors $\boldsymbol{U}$ and create error-prone exposures $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$. We fixed the validation study size at $n=100$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. Again, all estimates were empirically unbiased (Supplemental Figure S2).
The ETS-$PC_{1}^{*}$ offered the smallest total empirical coefficient variability across all models regardless of how severe the measurement error was (Figure 3B). The efficiency gap between the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs widened as $\sigma_{U}^{2}$ increased. Intuitively, the need for validation is greater when $\sigma^{2}_{U}$ is larger, such that $\boldsymbol{X}^{*}$ is noisier about $\boldsymbol{X}$. Meanwhile, the benefits of our proposed ETS-$PC_{1}^{*}$ design over ETS-$X_{1}^{*}$ and SRS grew as $\sigma_{U}^{2}$ increased. In other words, including information from all error-prone exposures, rather than just the primary one, seemed to become more important as the errors worsened.
The same trend was seen when breaking down the efficiency comparison by individual models, as well. As $\sigma_{U}^{2}$ increased, the efficiency of all designs and models decreased, and the advantages of the ETS-$PC_{1}^{*}$ design became more pronounced (Figure 4B). In fact, the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs offered almost the same efficiency in the primary model under the most severe setting where $\sigma_{U}^{2}=1$.
3.5 Varied Validation Proportion
We varied the proportion of the $N=1000$ patients whose exposures $\boldsymbol{X}$ were validated between $10\%$ (used elsewhere), $15\%$, and 25%. We kept $\sigma_{U}^{2}=1$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. The estimates were empirically unbiased with even just $10\%$ validated (Supplemental Figure S3).
For all validation proportions considered, the ETS-$PC_{1}^{*}$ design offered the smallest empirical total variability across the five models (Figure 3C). The gap between the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs was fairly consistent in magnitude across all three settings. As the validation proportion increased, fewer patients have missing data and need imputation, leading to better efficiency for all designs. Still, incorporating all five error-prone exposures through the ETS-$PC_{1}^{*}$ design remained beneficial. Regardless of how many patients were validated, the ETS-$PC_{1}^{*}$ design gained more information across all models, offering better statistical efficiency, than the ETS-$X_{1}^{*}$ and SRS.
When evaluating the models individually, the efficiency of all designs increased as more patients were validated, as expected (Figure 4C). The gains with the ETS-$PC_{1}^{*}$ design over ETS-$X_{1}^{*}$ and SRS were persistent and large for all validation proportions considered. This finding suggests that the ETS-$PC_{1}^{*}$ design could be beneficial for partial validation studies with a wide range of budgets and sizes.
3.6 Shared Outcome Model
We considered an alternative data generating mechanism where the five exposures $X_{1},\dots,X_{5}$ shared a common numeric outcome $Y$. Under this mechanism, the shared outcome was generated as $Y=\beta_{1}X_{1}+\beta_{2}X_{2}+\beta_{3}X_{3}+\beta_{4}X_{4}+\beta_{5}X_{5}+0.1Z+\epsilon$, where random errors $\epsilon$ come from a standard normal distribution. We fixed the validation study size $n=100$, error variance $\sigma_{U}^{2}=1$, and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as in Section 3.1. Three scenarios were included: (i) only $X_{1}$ has a relationship with $Y$ ($\beta_{1}=1$ and all other $\beta_{j}=0$), (ii) only $X_{2}$ has a relationship with $Y$ ($\beta_{2}=1$ and all other $\beta_{j}=0$), and (iii) all $\boldsymbol{X}$ have relationships with $Y$ (all $\beta_{1}=\cdots=\beta_{5}=0.2$). In all scenarios, separate linear models of $Y$ on $X_{j}$ and $Z$ were fit using single imputation. Through these settings, we compared the validation study designs when only one exposure was associated with $Y$ (and when it was/was not the one targeted with the ETS-$X_{1}^{*}$ design) versus when all exposures were associated.
Imputation continued to lead to unbiased estimates (Supplemental Figure S4), and an interesting new story emerges when comparing the designs’ efficiency. Essentially, ETS-$PC_{1}^{*}$ was the “safest” choice, since it offered the same if not better efficiency as ETS-$X_{1}^{*}$, both in terms of the total across all coefficients and in the individual models (Figure 5). As expected, when only $X_{1}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs offered nearly identical efficiency in the model using $X_{1}$. However, the ETS-$PC_{1}^{*}$ design offered noticeable efficiency gains over ETS-$X_{1}^{*}$ for the other models. When only $X_{2}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ design offered substantially better efficiency in total and for all five models separately, but especially for the model with $X_{2}$. Finally, when all $\boldsymbol{X}$ were associated with $Y$, the ETS-$PC_{1}^{*}$ design outperformed the ETS-$X_{1}^{*}$ and SRS designs for all models.
Figure 5: Simulation results comparing the empirical total coefficient variability across all models (A) and empirical efficiency for the exposure coefficient per model (B) under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. There was a shared outcome $Y$, and it was generated from the exposures $X_{1},\dots,X_{5}$ under scenarios where only one exposure was associated (only $\beta_{1}\neq 0$ for $X_{1}$ or only $\beta_{2}\neq 0$ for $X_{2}$) versus all covariates are associated (all $\beta_{j}\neq 0$ for $X_{1},\dots,X_{5}$).
4 Application to the Error-Prone Dietary Intake Exposures
Having established the relative advantages of our proposed ETS-$PC_{1}^{*}$ validation study design in synthetic data (Section 3), we further examined its performance in more realistic data. It was not feasible for us to conduct multiple validation studies under different design strategies in practice. Instead, we adopted a real-world dataset and added errors to it. Importantly, using real data, we no longer control the covariance between exposures, which heavily influences how best to select our validation subset. Comparing the statistical efficiency of $\widehat{{\boldsymbol{\beta}}}_{1}$ between same-size validation studies with different design strategies is our primary objective.
4.1 About the Survey and Data
The NHANES includes various demographic, examination, laboratory, and questionnaire data. We focus on the dietary interview portion, called the “What We Eat in America” (WWEIA), which captures individual-level daily consumption of $64$ nutrients and food components and is an incredibly valuable source for dietary intake data [16].
In the most recent WWEIA (2021202120232023), a representative sample of individuals of all ages was asked to complete two $24$-hour dietary intake recalls approximately $1$ week apart. The food components and quantities recorded were converted into dietary nutrients using the Food and Nutrient Database for Dietary Studies, which provides caloric and nutrient information [40]. The NHANES data set reports the cumulative intake of these nutrients for each recall.
4.2 Modeling Diet-Driven Health Outcomes
We considered five models of interest, each one linking an examination or laboratory outcome $Y_{j}$ to a corresponding dietary intake exposure $X_{j}$ (Table 1). Each model was adjusted by the individuals’ sex, age, race and ethnicity (Hispanic origin), and education level to account for possible confounding, denoted by $\mathbf{Z}$}. Each model followed the linear regression in (1), and estimating the exposure coefficients ${\boldsymbol{\beta}}_{1}$ across all models was our goal.
Outcome Dietary Intake Covariate Clinical Relevance
$Y_{1}:$ Vitamin D $X_{1}:$ Calcium Vitamin D facilitates calcium absorption [5].
$Y_{2}:$ Resting Heart Rate $X_{2}:$ Caffeine Caffeine can affect heart rate and cardiovascular function [39].
$Y_{3}:$ High-Density Lipoproteins (HDL) Cholesterol $X_{3}$: Saturated Fat Saturated fat intake increases HDL cholesterol [4].
$Y_{4}:$ Insulin $X_{4}$: Alcohol Alcohol consumption may influence insulin sensitivity [33].
$Y_{5}:$ Folate $X_{5}:$ Food Folate Low folate levels can lead to higher odds of anemia [27].
Table 1: The outcome of interest, associated dietary nutrient value, and the associated source data for the five models from the NHANES data. All exposures are sourced from the dietary interview portion of the National Health and Nutrition Examination Survey (NHANES), called “What We Eat in America.”
We focused on individuals who (i) completed the first telephone interview and (ii) were not missing any of the models’ outcomes, exposures, or confounders. A sample of $N=2388$ met these criteria and were used for analysis. While excluding individuals with missing data is not always advisable (e.g., when missing data patterns may be informative), we elected to do so here to simplify our illustration of the proposed validation study design.
4.3 Measurement Error in Dietary Intake
The NHANES dietary interview uses the Automated Multiple-Pass Method (AMPM) which is considered the most sophisticated methodology for 24-hour food recall [38]. ADPM begins with a list of all foods and beverages consumed, followed by probing questions, collecting the time and occasion of meals, detailed descriptions, and a final nudge for anything else consumed.
Two main limitations of the AMPM exist. First, it requires multiple days of recalls to obtain accurate mean measurements of intake, which is both costly and time-consuming. Second, it can induce self-report bias from psycho-social behaviors that result in under or over estimation of nutrient and energy intake, though underreporting is most common (see [38] and references therein). Weighed food records are the gold standard for recording portions and the most accurate measurements, but has a high burden on participants. [44]. Weighed food records require measuring all portions prior to consumption and any food waste that remains after a meal. A large-scale study such as the NHANES cannot implement weighed food records due to the size, cost, and burden. Measurement error may occur during the food recall and impact assessment on the relationship of dietary intake with other biomarkers.
We treat the available dietary intake exposures from the NHANES dataset as the error-free versions $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$. Calcium ($X_{1}$) and saturated fat ($X_{3}$) were moderately correlated with each other ($\hat{\rho}=0.59$) and weakly correlated with food folate ($X_{5}$) ($\hat{\rho}=0.38$ and $0.34$, respectively). Caffeine ($X_{2}$) and alcohol ($X_{4}$) exhibited minimal correlation with other exposures (all $\hat{\rho}\leq 0.16$). See Supplemental Figure S5A for the full correlation matrix.
We simulated error-prone $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*},X_{4}^{*},X_{5}^{*})$ by generating measurement errors $\boldsymbol{U}=(U_{1},U_{2},U_{3},U_{4},U_{5})$ from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\mathrm{diag}(77805.2,7630.9,70.5,119.1,4991.9)$, where $\mathrm{diag}(\cdot)$ denotes a diagonal matrix. The error variance was chosen to be $\sigma_{Uj}^{2}=\widehat{\textrm{Var}}(X_{j})/4$, i.e., $1/4$ times the estimated variance of the corresponding error-free exposure. Then, we combined added these errors to the NHANES data to get $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$. Supplemental Figure S6 contains scatterplots of $\boldsymbol{X}$ versus $\boldsymbol{X}^{*}$.
As anticipated, after adding measurement error, the correlations between the five dietary intake covariates all exposures decreased but persisted (Supplemental Figure S5B). For example, the correlation between $X_{1}^{*}$ (calcium) and $X_{3}^{*}$ (saturated fat) was $\hat{\rho}^{*}=0.46$, while that between their error-free versions was $\hat{\rho}=0.59$.
4.4 Partial Validation of Dietary Intake Covariates
All $N=2388$ individuals had outcomes $\boldsymbol{Y}=(Y_{1},Y_{2},Y_{3},Y_{4},Y_{5})$, error-prone dietary intake $\boldsymbol{X}^{*}$, and other confounders $\boldsymbol{Z}$ observed. However, only a subset of $n=250$ of them underwent “validation” and had error-free dietary intake $\boldsymbol{X}$ observed.
First, we included the SRS design, which, while theoretically valid, leads to the widest 95% confidence intervals ($95\%$ CIs) for ${\boldsymbol{\beta}}_{1}$. It leads to model estimates that are asymptotically consistent (i.e., converge in probability to the true coefficients) as both sample sizes ($N$ and $n$) go to infinity. However, there can be discrepancies between the SRS estimates and full validation in finite samples.
Second, we considered an ETS-$X_{1}^{*}$ design that focuses on the error-prone dietary intake exposure, $X_{1}^{*}$ (calcium). This approach might appeal if we were primarily interested in this model, so that we tailored our validation sampling strategy to focus on it. However, we may later decide that the other four models are also relevant, and then re-purpose the validation study data to estimate them. The ETS-$X_{1}^{*}$ design should offer the narrowest $95\%$ CI for Model 1, but it may not even outperform SRS for Models 2–5.
Third, we applied our proposed ETS-$PC_{1}^{*}$ design. The first principal component, $PC_{1}^{*}$, explained $35\%$ of the variability of $\boldsymbol{X}^{*}$. All error-prone nutrient intake exposures contributed to $PC_{1}^{*}$, with the largest loading of $0.59$ for $X_{1}^{*}$ (calcium) and the smallest of $0.14$ for $X_{4}^{*}$ (alcohol). For comparison, we calculated the average percent variability explained by $PC_{1}^{*}$ in simulations. When varying the covariance structures for $\boldsymbol{X}$, $28\%$ was explained under independence, $48\%$ under equal dependence, and $39\%$ under unequal dependence. As the exposure error variance $\sigma_{U}^{2}$ increased, the percent of variability explained by $PC_{1}^{*}$ dropped from $64\%$ to $48\%$. Based on the correlation matrix for the NHANES data (Supplemental Figure S5), their covariance structure resembled the unequal dependence scenario; the $35\%$ variability explained by $PC_{1}^{*}$ aligned well with the simulations, too. Plus, the magnitude of measurement errors in the nutrient intake exposures was fairly severe.
4.5 Estimates from Multiple Models
Figure 6A summarizes estimates from the five diet-driven health outcome models. First, we computed the “gold standard” model estimates using the error-free covariates $\boldsymbol{X}$ on all $N=2388$ individuals from the original NHANES data. The estimated slopes $\widehat{\beta}_{1j}$ for each of the five nutrient intake exposures was approximately 0. All other coefficient estimates are summarized in Table 2.
Figure 6: Coefficient estimates and corresponding $95\%$ confidence intervals (A) and relative widths of $95\%$ confidence intervals for $\beta_{1j}$, corresponding to each of the five nutrient intake exposures ($j\in\{1,\dots,5\}$) from the five diet-driven health outcome models (B). In addition to the gold standard analysis (with all individuals having validation data), we considered partial validation with simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on $PC_{1}^{*}$ (ETS-$PC_{1}^{*}$). The analyses with partial validation data used multiple imputation.
Then, we considered analyses with a validation study of $n=250$ individuals. For the remaining $2138$ unvalidated individuals, error-free nutrient intake exposures were multiply imputed $75$ times. In general, the three imputation analyses led to estimates for $\beta_{1j}$ that differed slightly from the gold standard values. However, they would become more similar if more validation data were available.
The total coefficient variability for the gold standard analysis was $\sum_{j=1}^{5}\widehat{V}(\hat{\beta}_{1j})=0.002$, while for the partially validated analyses using the SRS and ETS-$X_{1}^{*}$ designs it was =0.007=0.007. With the ETS-$PC_{1}^{*}$ design, the total variability of $0.004$ with the multiple imputation analysis was much closer to the gold standard. Additionally, the ETS-$PC_{1}^{*}$ design offered narrower $95\%$ CIs in all five models than the SRS or ETS-$X_{1}^{*}$ (Figure 6B). It had the largest advantages in Model $4$, which was interesting given that the fourth nutrient intake exposure $X_{4}^{*}$ (alcohol) was not very correlated and had the smallest loading on $PC_{1}^{*}$. The ETS-$X_{1}^{*}$ design did not offer much improvement over SRS. It was particularly surprising for ETS-$X_{1}^{*}$ to not perform best for the corresponding model with $X_{1}$, as it did in simulations. However, the error-prone nutrient intake exposure could be so noisy that sampling on the extremes of $X_{1}^{*}$ did not necessarily lead us to validate the extremes of $X_{1}$ (which are the source of the efficiency gain).
5 Discussion
In simulations, our proposed sampling strategy improved the overall efficiency of the two-phase estimates across multiple models of interest under mild assumptions. The improvement was evident under varying covariance structures, measurement error severities, and validation proportions. Models where the error-prone exposure was strongly correlated with the others saw the greatest improvement. Still, when the exposures’ variability was relatively small compared to the outcomes’, any ETS procedure will see decreased benefits as the improved precision of exposures will not substantially improve the precision of the overall effect estimates.
We demonstrate the ETS-$PC_{1}^{*}$ design’s advantages in single and multiple imputation analyses, where model specification is critical to achieving unbiased estimates. In practice, selection procedures (e.g., based on information criteria) can help select the “best-fitting” imputation model [32]. However, the advantages of sampling on the first principal component should also manifest if applied to other model-based estimation procedures [20], which opens up possibilities for more robust analytical approaches, like semiparametric MLE [36].
Toward broader applicability of the proposed sampling strategy, there are multiple interesting directions for future work. First, investigations into efficiency gains with other common outcome types (including binary, count, or time-to-event) are needed. Second, adopting other existing designs for a dimension-reduced version of the error-prone exposures (like the first principal component) could offer further advantages. Third, modifications to other error settings (e.g., error-prone outcomes) are promising; in particular, the set of variables contributing to the PCA will need to be revisited.
6 Conclusion
Extreme tail sampling on the first principal component offers an accessible, intuitive way to balance multiple, competing analytical priorities when choosing which individuals to validate in studies with exposure measurement error. This strategy amplifies the benefits of a conventional single-objective two-phase design by avoiding the expenses of conducting separate, targeted validation samples for distinct models. It does so without incurring computational complexity, and we provide the user-friendly R auditDesignR package at https://github.com/sarahlotspeich/auditDesignR.
Supplementary Materials
Coefficient Gold Standard SRS $\boldsymbol{X_{1}^{*}} ETS$-$\boldsymbol{PC_{1}^{*}}$
Model 1: Vitamin D ($Y_{1}$) and Calcium ($X_{1}$)
Intercept $21.4$ ($11.9$, $30.9$) $23.3$ ($12.9$, $33.7$) $18.6$ ($7.9$, $29.3$) $20.2$ ($10.2$, $30.2$)
Nutrient Intake 0 (0, 0) 0 (0, 0) 0 (0, $0.01$) 0 (0, $0.01$)
Female $12.3$ ($9.5$, $15.1$) $11.9$ ($8.9$, $15$) $12.6$ ($9.7$, $15.5$) $12.3$ ($9.5$, $15.1$)
Age $0.7$ ($0.6$, $0.8$) $0.7$ ($0.6$, $0.8$) $0.7$ ($0.6$, $0.8$) $0.7$ ($0.6$, $0.8$)
Race and Ethnicity
   Other Hispanic 44 (2.7-2.7, $10.8$) $3.9$ (2.9-2.9, $10.7$) $4.3$ (2.5-2.5, $11.1$) $4.4$ (2.4-2.4, $11.3$)
   Non-Hispanic White $14.6$ ($8.8$, $20.4$) $14.7$ ($8.9$, $20.6$) $14.5$ ($8.6$, $20.3$) $14.6$ ($8.8$, $20.4$)
   Non-Hispanic Black 4.2-4.2 (11.1-11.1, $2.7$) 4.4-4.4 (11.3-11.3, $2.6$) 4.3-4.3 (11.2-11.2, $2.6$) 4.2-4.2 (11.1-11.1, $2.7$)
   Other Race 88 ($0.9$, $15$) $7.9$ ($0.9$, $14.9$) $7.9$ ($0.8$, $14.9$) $7.8$ ($0.7$, $14.8$)
Education Level
   9119-11th Grade 7.9-7.9 (16.8-16.8, $1.1$) 7.8-7.8 (16.8-16.8, $1.1$) 8.1-8.1 (17.1-17.1, $0.8$) 7.8-7.8 (16.7-16.7, $1.2$)
   High School Grad/GED or Equivalent $0.1$ (7.7-7.7, $7.9$) $0.1$ (7.7-7.7, $7.9$) $0.2$ (7.7-7.7, 88) $0.2$ (7.6-7.6, 88)
   Some College or AA Degree $4.7$ (3-3, $12.4$) $4.8$ (2.9-2.9, $12.5$) $4.6$ (3.1-3.1, $12.3$) $4.6$ (3.1-3.1, $12.3$)
   College Graduate or Above $7.4$ (0.3-0.3, $15$) $7.5$ (0.1-0.1, $15.2$) $7.3$ (0.3-0.3, $14.9$) $7.3$ (0.3-0.3, $14.9$)
Model 2: Resting Heart Beat ($Y_{2}$) and Caffeine ($X_{2}$)
Intercept $72$ ($68.8$, $75.3$) $71.6$ ($68.4$, $74.8$) $72$ ($68.7$, $75.2$) $71.9$ ($68.7$, $75.1$)
Nutrient Intake 0 (0, 0) 0 (0, $0.01$) 0 (0.01-0.01, 0) 0 (0, 0)
Female $3.1$ ($2.2$, $4.1$) $3.3$ ($2.4$, $4.3$) $3.2$ ($2.2$, $4.1$) $3.2$ ($2.2$, $4.2$)
Age 0.1-0.1 (0.1-0.1, 0.1-0.1) 0.1-0.1 (0.2-0.2, 0.1-0.1) 0.1-0.1 (0.2-0.2, 0.1-0.1) 0.1-0.1 (0.2-0.2, 0.1-0.1)
Race and Ethnicity
   Other Hispanic $3.2$ ($0.9$, $5.6$) $3.3$ ($0.9$, $5.7$) $3.2$ ($0.9$, $5.6$) $3.3$ ($0.9$, $5.6$)
   Non-Hispanic White $2.8$ ($0.8$, $4.9$) $2.4$ ($0.4$, $4.5$) $2.8$ ($0.7$, $4.9$) $2.8$ ($0.7$, $4.8$)
   Non-Hispanic Black $1.8$ (0.6-0.6, $4.2$) $2.2$ (0.3-0.3, $4.6$) $1.9$ (0.6-0.6, $4.3$) $1.9$ (0.5-0.5, $4.3$)
   Other Race $3.5$ (11, $5.9$) $3.2$ ($0.7$, $5.6$) $3.5$ (11, 66) $3.4$ (11, $5.9$)
Education Level
   9119-11th Grade 22 (1.2-1.2, $5.1$) $1.4$ (1.8-1.8, $4.6$) 22 (1.2-1.2, $5.1$) $1.9$ (1.2-1.2, $5.1$)
   High School Grad/GED or Equivalent $1.3$ (1.4-1.4, 44) $1.1$ (1.6-1.6, $3.9$) $1.3$ (1.4-1.4, 44) $1.3$ (1.4-1.4, 44)
   Some College or AA Degree $1.4$ (1.2-1.2, $4.1$) 11 (1.7-1.7, $3.7$) $1.4$ (1.3-1.3, $4.1$) $1.4$ (1.3-1.3, $4.1$)
   College Graduate or Above 1.3-1.3 (3.9-3.9, $1.4$) 1.6-1.6 (4.3-4.3, $1.1$) 1.3-1.3 (3.9-3.9, $1.4$) 1.3-1.3 (4-4, $1.4$)
Model 3: HDL Cholesterol ($Y_{3}$) and Saturated Fat ($X_{3}$)
Intercept $36.5$ ($32.6$, $40.5$) $36.3$ ($32.1$, $40.4$) $36.4$ ($32.2$, $40.6$) $35.8$ ($31.6$, $40$)
Nutrient Intake $0.01$ (0.03-0.03, $0.04$) $0.02$ (0.04-0.04, $0.08$) $0.01$ (0.05-0.05, $0.07$) $0.03$ (0.02-0.02, $0.08$)
Female $8.8$ ($7.7$, $10$) $8.9$ ($7.7$, $10.1$) $8.8$ ($7.7$, $10$) $8.9$ ($7.8$, $10.1$)
Age $0.1$ ($0.1$, $0.1$) $0.1$ ($0.1$, $0.1$) $0.1$ ($0.1$, $0.1$) $0.1$ ($0.1$, $0.1$)
Race and Ethnicity
   Other Hispanic $1.6$ (1.2-1.2, $4.4$) $1.5$ (1.3-1.3, $4.3$) $1.6$ (1.2-1.2, $4.4$) $1.7$ (1.1-1.1, $4.5$)
   Non-Hispanic White $3.6$ ($1.2$, 66) $3.5$ (11, 66) $3.6$ ($1.2$, 66) $3.5$ ($1.1$, $5.9$)
   Non-Hispanic Black $6.4$ ($3.5$, $9.2$) $6.3$ ($3.4$, $9.2$) $6.4$ ($3.5$, $9.2$) $6.4$ ($3.5$, $9.2$)
   Other Race $3.5$ ($0.5$, $6.4$) $3.3$ ($0.3$, $6.3$) $3.4$ ($0.5$, $6.3$) $3.3$ ($0.4$, $6.3$)
Education Level
   9119-11th Grade $1.4$ (2.3-2.3, $5.1$) $1.4$ (2.4-2.4, $5.1$) $1.4$ (2.3-2.3, $5.1$) $1.4$ (2.3-2.3, $5.1$)
   High School Grad/GED or Equivalent $2.9$ (0.3-0.3, $6.2$) $2.9$ (0.3-0.3, $6.2$) $2.9$ (0.3-0.3, $6.2$) $2.9$ (0.3-0.3, $6.2$)
   Some College or AA Degree $3.1$ (0.1-0.1, $6.3$) $3.1$ (0.1-0.1, $6.3$) $3.1$ (0.1-0.1, $6.3$) $3.1$ (0.1-0.1, $6.3$)
   College Graduate or Above $6.4$ ($3.3$, $9.6$) $6.4$ ($3.3$, $9.6$) $6.4$ ($3.2$, $9.6$) $6.4$ ($3.3$, $9.6$)
Table 2: Estimates from all fitted models from the application to error-prone dietary intake exposures in the National Health and Nutrition Examination Survey (NHANES) data. The Gold Standard estimates used only the error-free $\boldsymbol{X}$ from NHANES for all individuals. All others used $\boldsymbol{X}$ from a subset of $n=250$ individuals and multiply imputed them from $\boldsymbol{X}^{*}$ and $\mathbf{Z}$} for the rest. Three different validation study designs were considered: simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$).
Coefficient Gold Standard SRS $\boldsymbol{X_{1}^{*}} ETS$-$\boldsymbol{PC_{1}^{*}}$
Model 4: Insulin ($Y_{4}$) and Alcohol ($X_{4}$)
(Intercept) $18.5$ ($12.9$, $24.1$) $18.6$ ($13$, $24.3$) $18.6$ ($13$, $24.2$) $18.9$ ($13.3$, $24.6$)
Nutrient Intake 0.03-0.03 (0.07-0.07, $0.01$) 0.06-0.06 (0.15-0.15, $0.03$) 0.08-0.08 (0.18-0.18, $0.02$) 0.07-0.07 (0.12-0.12, 0.02-0.02)
Female 2.6-2.6 (4.3-4.3, 0.9-0.9) 2.7-2.7 (4.4-4.4, 0.9-0.9) 2.8-2.8 (4.6-4.6, 1.1-1.1) 2.9-2.9 (4.6-4.6, 1.1-1.1)
Age 0 (0.1-0.1, 0) 0 (0.1-0.1, 0) 0 (0.1-0.1, 0) 0 (0.1-0.1, 0)
Race and Ethnicity
   Other Hispanic 0.2-0.2 (4.3-4.3, $3.9$) 0.3-0.3 (4.4-4.4, $3.9$) 0.1-0.1 (4.2-4.2, $4.1$) 0.3-0.3 (4.5-4.5, $3.8$)
   Non-Hispanic White 3.4-3.4 (7-7, $0.2$) 3.3-3.3 (6.9-6.9, $0.3$) 3.2-3.2 (6.8-6.8, $0.3$) 3.3-3.3 (6.9-6.9, $0.2$)
   Non-Hispanic Black 3-3 (7.2-7.2, $1.3$) 2.7-2.7 (7-7, $1.6$) 2.8-2.8 (7-7, $1.5$) 2.8-2.8 (7.1-7.1, $1.4$)
   Other Race 4.4-4.4 (8.7-8.7, 0.1-0.1) 4.4-4.4 (8.7-8.7, 0.1-0.1) 4.3-4.3 (8.6-8.6, 0) 4.4-4.4 (8.7-8.7, 0.1-0.1)
Education Level
   9119-11th Grade $1.7$ (3.8-3.8, $7.2$) $1.9$ (3.6-3.6, $7.4$) $2.3$ (3.2-3.2, $7.9$) $2.1$ (3.4-3.4, $7.6$)
   High School Grad/GED or Equivalent $1.8$ (3-3, $6.6$) $1.8$ (3-3, $6.6$) 22 (2.8-2.8, $6.8$) $1.7$ (3-3, $6.5$)
   Some College or AA Degree $0.9$ (3.8-3.8, $5.6$) 11 (3.7-3.7, $5.7$) $1.1$ (3.6-3.6, $5.8$) $1.1$ (3.6-3.6, $5.8$)
   College Graduate or Above 1.3-1.3 (5.9-5.9, $3.4$) 1-1 (5.7-5.7, $3.7$) 1.1-1.1 (5.8-5.8, $3.6$) 1-1 (5.7-5.7, $3.6$)
Model 5: Folate ($Y_{5}$) and Food Folate ($X_{5}$)
(Intercept) $312.2$ ($246.9$, $377.5$) $279.6$ ($208.5$, $350.7$) $334.5$ ($262.3$, $406.7$) $319.1$ ($249.4$, $388.8$)
Nutrient Intake 0.01-0.01 (0.08-0.08, $0.05$) $0.11$ (0.01-0.01, $0.23$) 0.09-0.09 (0.2-0.2, $0.03$) 0.04-0.04 (0.14-0.14, $0.07$)
Female $26$ ($6.9$, $45.1$) $32.4$ ($12.4$, $52.5$) $22.2$ ($2.4$, $42$) $24.7$ (55, $44.4$)
Age $3.5$ ($2.9$, 44) $3.5$ ($2.9$, $4.1$) $3.4$ ($2.8$, 44) $3.4$ ($2.9$, 44)
Race and Ethnicity
   Other Hispanic 15.2-15.2 (61.6-61.6, $31.2$) 7.9-7.9 (55-55, $39.2$) 18.1-18.1 (64.7-64.7, $28.5$) 15.5-15.5 (62-62, $30.9$)
   Non-Hispanic White $48$ ($8.2$, $87.9$) $52.1$ ($12$, $92.2$) $45.2$ ($5.2$, $85.3$) $47.2$ ($7.3$, $87.2$)
   Non-Hispanic Black 57-57 (104.4-104.4, 9.5-9.5) 51.2-51.2 (99.1-99.1, 3.2-3.2) 58.9-58.9 (106.5-106.5, 11.4-11.4) 57.2-57.2 (104.7-104.7, 9.7-9.7)
   Other Race 7.6-7.6 (55.8-55.8, $40.6$) 4.4-4.4 (52.8-52.8, $44.1$) 5.3-5.3 (53.7-53.7, $43.1$) 6.6-6.6 (54.9-54.9, $41.7$)
Education Level
   9119-11th Grade 25.8-25.8 (87.1-87.1, $35.5$) 29.5-29.5 (91.1-91.1, $32.2$) 26-26 (87.3-87.3, $35.3$) 25.6-25.6 (86.9-86.9, $35.8$)
   High School Grad/GED or Equivalent $1.2$ (52.3-52.3, $54.8$) 0.2-0.2 (53.8-53.8, $53.4$) 1.8-1.8 (55.5-55.5, $51.9$) $1.1$ (52.5-52.5, $54.6$)
   Some College or AA Degree $25.2$ (27.5-27.5, $77.9$) $24.6$ (28.2-28.2, $77.4$) $22.2$ (30.7-30.7, $75$) $25.3$ (27.5-27.5, $78$)
   College Graduate or Above $19.5$ (32.8-32.8, $71.8$) $10.8$ (42.3-42.3, $63.9$) $20.8$ (31.5-31.5, $73$) $20.2$ (32.2-32.2, $72.5$)
Table 2 (Continued): Estimates from all fitted models from the application to error-prone dietary intake exposures in the National Health and Nutrition Examination Survey (NHANES) data. The Gold Standard estimates used only the error-free $\boldsymbol{X}$ from NHANES for all individuals. All others used $\boldsymbol{X}$ from a subset of $n=250$ individuals and multiply imputed them from $\boldsymbol{X}^{*}$ and $\mathbf{Z}$} for the rest. Three different validation study designs were considered: simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$).